# X-GRM: Large Gaussian Reconstruction Model for Sparse-view X-rays to Computed Tomography

### v1

@@ 5 Conclusion
This paper presents X-GRM, a large feedforward model for reconstructing 3D CT volumes from sparse-view X-rays using a scalable transformer architecture and Voxel-based Gaussian Splatting representation. Trained on the collected large ReconX-15K dataset, X-GRM demonstrates superior reconstruction quality with varied number of X-ray inputs. By addressing previous limitations in model capacity, volume representation, and training data scale, our work represents an advancement in non-invasive sparse-view CT imaging for clinical applications.
@@ A.2 Dataset details
As stated in the main text, we collect 8 public datasets covering organs and structures including chest, abdomen, pelvis, and teeth. We apply data standardization to adjust the data through three steps: resampling spacing, adjusting resolution to 2563, clipping HU values, and normalizing to $[0,1]$. Then, we split each dataset into training/validation/test splits with ratio 20/1/1. Meanwhile, as most datasets solely contain CT volumes, we utilize the TIGRE toolbox [3] to render X-ray projections. The specific parameter settings for each dataset are shown in the Tab. 7.
Table 7: The parameters used to pre-process each dataset of ReconX-15K.
@@ A.5 Broader impacts
Impacts on research community. The ReconX-15K dataset establishes a comprehensive benchmark for CT reconstruction methods, enabling standardized evaluation and fair comparison. This work demonstrates the transformative potential of large-scale data and high-capacity models in medical imaging, encouraging similar approaches in related domains. By open-sourcing code, models, and data, X-GRM contributes essential resources for advancing medical image reconstruction research.

### v2

@@ 5 Conclusion
This paper presents X-GRM, a large feedforward model for reconstructing 3D CT volumes from sparse-view X-rays using a scalable transformer architecture and Voxel-based Gaussian Splatting representation. Trained on a large CT reconstruction dataset, X-GRM demonstrates superior reconstruction quality with sparse-view X-ray inputs. By addressing previous limitations in model capacity and volume representation, our work represents an advancement in non-invasive sparse-view CT imaging for clinical applications.
@@ A.2 Dataset details
As stated in the main text, we follow X-LRM [77] to collect 8 public datasets covering organs and structures including chest, abdomen, pelvis, and teeth. We apply data standardization to adjust the data through three steps: resampling spacing, adjusting resolution to 2563, clipping HU values, and normalizing to $[0,1]$. Then, we split each dataset into training/validation/test splits with ratio 20/1/1. Meanwhile, as most datasets solely contain CT volumes, we utilize the TIGRE toolbox [3] to render X-ray projections. The specific parameter settings for each dataset are shown in the Tab. 7.
Table 7: The parameters used to pre-process each subset of the collected dataset.
@@ A.5 Broader impacts
Impacts on research community. This work demonstrates the transformative potential of large-scale data and high-capacity models in medical imaging, encouraging similar approaches in related domains. By open-sourcing code, models, and data, X-GRM contributes essential resources for advancing medical image reconstruction research.